OpenStudy (anonymous):

Find the sum of the measures of the angles in a pentagon. a. 360° b. 540° c. 620° d. 900°

OpenStudy (mrnood):

 Sum of external angles is 360 deg (there are 5 of them in a pentagon...) e+i = 180 for each side...

OpenStudy (acxbox22):

180(n-2)= sum of the measures of the angles where n is the number of sides

OpenStudy (mrnood):

@acxbox22 it is discouraged to give complete answers. The OP had all the info in my answer to deduce the formula you gave.

OpenStudy (acxbox22):

@MrNood you gave the user the answer for the number of external sum of angles i gave them the formula for the sum of the internal sum of angles

OpenStudy (mrnood):

no I didn't - I said e+i = 180 THAT is enough to deduce the answer

OpenStudy (acxbox22):

and in the CoC it says it is discouraged to give direct answers not explanations and answers

OpenStudy (mrnood):

Don't devalue the question/answer process! •Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. •Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in! Give Help, Not Answers –I will encourage and guide those needing help, and not just give them an answer

OpenStudy (acxbox22):

i did not "just give them an answer"

OpenStudy (acxbox22):

here to make you happy...i will delete my post \(\Huge\color{green}{\ddot\smile}\)

OpenStudy (mrnood):

Giving the formula wouyld suffice if you had to - there was no need ot work out th enumbers for them. 180(5-2)=180(3)=540 degrees = ANSWER

OpenStudy (mrnood):

It's not to make me happy - I couldn't care less. The OP learns NOTHING from copying your answer - especially when the question is a lame multichoice option.